OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method (1 point) Part 3: Provide the solution to the system. (2 points) x – 3y + 2z = –12 x + 2y + 3z = 6 2x – 3y – z = –2

OpenStudy (mertsj):

Part 1: I would use elimination. Part 2: I chose that method because I think it is the easiest. Part 3: Subtract equation 2 from equation 1: -5y-z=-18 Multiply equation 2 by -2: -2x-4y-6z=-12 Add equation 3: 2x-3y-z=-2 -7y-7z=-14 Divide that equation by -7: y+z=2 Add -5y-z=-18 -4y=-16 y = 4 Substitute in y +z=2 4+z=2 z=-2 Substitute y and z in the second equation: x+2(4)+3(-2)=6 x=4 Solution: (4,4,-2)

OpenStudy (anonymous):

You're amazing